# RAGSys: Item-Cold-Start Recommender as RAG System

### v1

@@ 3. Methodology
3.1. Problem Statement
Consider an answer $a$ that should be generated by an LLM in response to a query $q$. The query can be a simple question such as ”Who was the first man to walk on the moon?”, or more general message such as ”I’d like to find red shoes”. The answer could take on many forms, such a factual response ”Neil Armstrong”, a clarifying question ”What size and style of shoe are you looking for?”, a JSON payload to send to an API ”{"search_terms":["red","shoe"]}” etc.
@@ 3.2.1. Classification Metrics
We also define a metric MC2 which extends this further to the case that multiple correct answers $\mathbf{a}$ and multiple incorrect answers $\mathbf{\bar{a}}$ are provided for each query. This metric is the the average number of correct answers which have greater probability than all incorrect answers.
$\frac{1}{|\mathcal{D}|}\sum_{(q,\mathbf{a},\mathbf{\bar{a}})\in\mathcal{D}}\frac{1}{|\mathbf{a}|}\sum_{a\in\mathbf{a}}\prod_{\bar{a}\in\mathbf{\bar{a}}}\mathbf{1}\Big{(}p_{M}\big{(}a\mid q,R(q)\big{)}>p_{M}\big{(}\bar{a}\mid q,R(q)\big{)}\Big{)}$
@@
$\frac{1}{|\mathcal{D}|}\sum_{(q,\mathbf{a},\mathbf{\bar{a}})\in\mathcal{D}}\frac{\sum_{a\in\mathbf{a}}p_{M}\big{(}a\mid q,R(q)\big{)}}{\sum_{\bar{a}\in\mathbf{\bar{a}}}p_{M}\big{(}\bar{a}\mid q,R(q)\big{)}}$
These metrics and their names follow those defined in (Lin et al., 2021). While they are easy to interpret, these metrics are not well normalized: they don’t take into account all possible correct and incorrect answers. As a result, if the sample of correct and incorrect answers have varying lengths and use of rare vocabulary tokens, these will impact the metrics.
3.2.2. Direct Preference Optimization Metric
@@
We define the DPO metric as the negative of the Direct Preference Optimization loss (Rafailov et al., 2023), which satisfies these properties:
$\log\sigma\Big{(}\log\frac{p_{M}\big{(}a\mid q,R(q)\big{)}}{p_{M}(a\mid q)}-\log\frac{p_{M}\big{(}\bar{a}\mid q,R(q)\big{)}}{p_{M}(\bar{a}\mid q)}\Big{)}$
@@ 3.3. Retrieval Algorithm
3.3.1. Query Relevance
The relevance score between the query and each candidate demonstration is calculated using the cosine similarity of their respective BERT embeddings (Devlin et al., 2019). By computing the cosine similarity between the query embedding and the embedding of each demonstration’s query, we obtain a score that reflects the topical similarity and semantic alignment between the query and the candidate demonstration.
3.3.2. Retrieved Demonstrations Diversity
@@
3.3.3. Demonstration Quality Bias
@@
This can be interpreted as measuring the probability of the correct answer $a$ given the query $q$, normalized to the length of the answer. This can also been interpreted as a proxy for a popularity bias, akin to the number of connections of an item in graph-based retrieval algorithms like recommender systems. Like in the article (Gonen et al., 2023), the intuition is that the more frequently a related sequence of tokens occurs in the pre-training dataset of the LLM, the more likely the model will be able to extract its relevant information. Rather than directly analyzing the massive amount of text data (often trillions of tokens) used to pre-train the LLM, we focus on the perplexity of the sequence. Perplexity acts as a proxy, indicating how surprised the LLM is by the sequence, essentially, how well it aligns with what the LLM expects to see. A parameter $\lambda_{b}$ controls the trade-off between relevance/diversity and quality bias. Lower values of $\lambda_{b}$ emphasize high-quality demonstrations.
@@
4. Experiments
4.1. Experimental Setup
4.1.1. The Dataset Choice
@@
4.1.2. The TruthfulQA Dataset
We chose to conduct our experiments using the TruthfulQA dataset (Lin et al., 2021) which meets these requirements. The dataset contains 817 distinct examples, which yields a standard error in the range of 1% to 2% for accuracy measures in the range of 90% to 50%. Each example contains a single query, and a variable number of correct and incorrect answers. And by considering each distinct $(q,a)$ pair as a demonstration for the purpose of building a context, the retriever is faced with similar demonstrations as multiple $(q,a)$ pairs share the same query (on average, the dataset contains 3.5 correct answers for each query).
4.1.3. Generating Demonstrations Pairs and Triplets
@@
  • $\mathcal{D}_{MC}$: this is the dataset as described in (Lin et al., 2021). It contains 817 examples, each of which contains a variable number of correct and incorrect answers. The metrics MC1, MC2 and MC3, which can accept an arbitrary number of correct and incorrect answers as inputs, are calculated over this dataset.
@@
4.1.4. The Language Models
We conducted our experiments using two noteworthy LLMs: Mistral-7B-v0.1 (Jiang et al., 2023) and Llama-3-8B-chat  [^1] https://llama.meta.com/llama3/. Both models are open-weights LLMs, meaning their internal parameters are publicly available for scrutiny and potential fine-tuning. These modern models stand out for achieving impressive performance on various tasks despite their relatively compact size. This efficiency makes it an attractive option for resource-constrained environments where deploying colossal models might not be feasible.
4.2. Implementation Details and Ablation Study
We implemented the Algorithm 1 and metrics from Section 3.2.1 in python and ran the experiments using an external LLM API. We did not perform hyper-parameter tuning, and fixed the two parameters to $\lambda_{d}=0.75$ and $\lambda_{b}=0.95$ in all experiments. We fixed the amount of retrieved demonstrations to $k=6$, matching the number of few-shot examples from the fixed primer example from the TruthfulQA paper (Lin et al., 2021).
@@ 4.2. Implementation Details and Ablation Study
  • Fix: fixed primer examples (Lin et al., 2021)
  • Bias: Pure quality bias (Gonen et al., 2023)
  • Rel: Pure semantic similarity (Liu et al., 2022a) (KATE)
@@
  • Rel+Div: Semantic similarity plus diversity (Ye et al., 2023a)
@@
4.3. Main Results
We present the experimental metrics for the 6 retrievers in Table 1 for Mistral-7B-v0.1 and Table 2 for Llama-3-8B.
Table 1. Evaluation Metrics with Mistral-7B-v0.1
@@ 4.3. Main Results
Table 2. Evaluation Metrics with Llama-3-8B-chat
@@
The results consistently demonstrate that incorporating both relevance and diversity into the retrieval strategy leads to superior performance across all metrics and for both LLMs. This finding reinforces the importance of not just retrieving relevant demonstrations but also ensuring a diverse set that maximizes the informative value for the LLM.
Interestingly, the impact of the low perplexity bias yields contrasting results. For Mistral-7B, adding this bias results in a decline in performance on almost all metrics. Conversely, Llama-3-8B exhibits overall improvement with the low perplexity bias. This intriguing observation suggests that LLM-dependent hyper-parameter tuning of $\lambda_{b}$ might be necessary to optimize retrieval strategies for specific models. Alternatively, the low perplexity bias itself may benefit from further refinement. Using an opposite intuition, we may argue that instead of prioritizing demonstrations the LLM already finds likely, introducing demonstrations that surprise the model the most could be beneficial for certain LLMs, potentially maximizing the learning impact of each demonstration. These findings open exciting new avenues for future research in ICL retrieval strategies, creating a parallel with novelty and serendipity concepts in recommender systems.
4.4. Calibrating Diversity using DPO
@@
Figure 1. Diversity Metric and DPO. Non-monotonous relationship between a diversity metric, the average cosine similarity between embedding pairs, and the quality metric DPO. Obtained by variying $\lambda_{d}$ with Llama-3-8B-chat and $k=6$.
[…]
The relationship between diversity metric and DPO is a noisy concave function.
5. Discussion: Real-World RAG Systems
@@
5.1. Balancing Performance and Efficiency
Retrieval latency is crucial at scale. Exhaustive, brute-force nearest neighbor search is computationally expensive and impractical. Instead, real-world systems leverage efficient indexing techniques and approximated kNN algorithms, as described in (Beygelzimer et al., 2006), to ensure fast retrieval times. This approach is essential for handling large datasets while maintaining responsiveness. To seamlessly integrate with existing retrieval engines and leverage their optimized search capabilities, a retrieval algorithm for RAG must ensure its data is stored in a format compatible with these engines. Commonly indexed data structures include text itself or low-dimensional dense vector embeddings. By adhering to these indexing practices, RAG systems can effectively leverage the power of existing retrieval engines and achieve fast, scalable retrieval of informative examples.
5.2. ICL RAG versus Fine-Tuning
@@
An interesting contrast between ICL and fine-tuning is highlighted in the paper (Liu et al., 2022b). The paper argues that fine-tuning can be more efficient than few-shot ICL in terms of cost and latency due to the super-linear increase in LLM latency with growing prompt sizes. However, this latency concern is less relevant with inference-throughput optimized LLM systems built with large GPU clusters, such as commonly used APIs. In these systems, the observed latency remains independent of the prompt size. From a latency perspective, adding ICL demonstrations can be considered free. Additionally, the paper suggests that ICL requires scanning through 100% of the demonstrations at query time. However this does not hold when employing real retrieval engines with indexing and approximate kNN, which significantly reduce the number of examples scanned during retrieval.
@@
5.3. Achieving State-of-the-Art Retrieval with Available Tools
@@
Due to the complexities of incorporating such rules directly within retrieval indexing algorithm (Li and Chan, 2013), none of the solutions known to the authors from any of the above category implements diversity or quality biasing of the result. A common heuristic to mitigate this problem is to retrieve a larger set of candidate examples (e.g., double the desired number) and then apply diversity techniques like MMR on the retrieved candidates as a post-processing step. Quality biasing can be indirectly achieved by modifying the indexed embeddings themselves. For instance, reducing the norm of embeddings associated with low-quality content can nudge the retrieval algorithm towards higher-quality examples. An exact implementation in the context of cosine-similarity or dot-product relevance is to add an additional column storing the quality bias, and set the corresponding value to $1$ in the embedding of the query.
While vector search offers a powerful foundation for practical ICL retrieval, it often lacks native support for essential considerations like diversity or quality bias. These aspects are crucial for ensuring informative and effective retrieval of few-shot learning examples. Existing tools for recommendation engines, on the other hand, often excel in these areas. Recommendation engines natively incorporate rules that promote diversity by recommending a variety of items, or quality bias by prioritizing most popular products. Future research directions as well as practical systems for ICL retrieval could explore adapting or integrating these well-established techniques from recommender systems to further enhance the effectiveness and sophistication of few-shot learning through information retrieval. State-of-the-art ICL for real-world applications can be achieved by combining the strengths of vector search with the established ”diversity-aware” retrieval approaches from recommender systems.
5.4. Conclusion

### v2

@@ 3 Methodology
3.1 Problem Statement
Consider an answer $a$ that should be generated by an LLM in response to a query $q$. The query can be a simple question such as "Who was the first man to walk on the moon?", or more general message such as "I’d like to find red shoes". The answer could take on many forms, such a factual response "Neil Armstrong", a clarifying question "What size and style of shoe are you looking for?", a JSON payload to send to an API "{"search_terms":["red","shoe"]}" etc.
@@ 3.2.1 Classification Metrics
We also define a metric MC2 which extends this further to the case that multiple correct answers $\mathbf{a}$ and multiple incorrect answers $\mathbf{\bar{a}}$ are provided for each query. This metric is the average number of correct answers which have greater probability than all incorrect answers.
[…]
@@
[…]
These metrics and their names follow those defined in [48]. While they are easy to interpret, these metrics are not well normalized: they don’t take into account all possible correct and incorrect answers. As a result, if the sample of correct and incorrect answers have varying lengths and use of rare vocabulary tokens, these will impact the metrics.
3.2.2 Direct Preference Optimization Metric
@@
We define the DPO metric as the negative of the Direct Preference Optimization loss [49], which satisfies these properties:
$\text{DPO}=\log\sigma\Big{(}\log\frac{p_{M}\big{(}a\mid q,R(q)\big{)}}{p_{M}(a\mid q)}-\log\frac{p_{M}\big{(}\bar{a}\mid q,R(q)\big{)}}{p_{M}(\bar{a}\mid q)}\Big{)}$
@@ 3.3 Retrieval Algorithm
3.3.1 Query Relevance
The relevance score between the query and each candidate demonstration is calculated using the cosine similarity of their respective BERT embeddings [21]. By computing the cosine similarity between the query embedding and the embedding of each demonstration’s query, we obtain a score that reflects the topical similarity and semantic alignment between the query and the candidate demonstration.
3.3.2 Retrieved Demonstrations Diversity
@@
3.3.3 Demonstration Quality Bias
@@
This can be interpreted as measuring the probability of the correct answer $a$ given the query $q$, normalized to the length of the answer. This can also been interpreted as a proxy for a popularity bias, akin to the number of connections of an item in graph-based retrieval algorithms like recommender systems. Like in the article [12], the intuition is that the more frequently a related sequence of tokens occurs in the pre-training dataset of the LLM, the more likely the model will be able to extract its relevant information. Rather than directly analyzing the massive amount of text data (often trillions of tokens) used to pre-train the LLM, we focus on the perplexity of the sequence. Perplexity acts as a proxy, indicating how surprised the LLM is by the sequence, essentially, how well it aligns with what the LLM expects to see. A parameter $\lambda_{b}$ controls the trade-off between relevance/diversity and quality bias. Lower values of $\lambda_{b}$ emphasize high-quality demonstrations.
@@
4 Experiments
4.1 Experimental Setup
4.1.1 The Dataset Choice
@@
4.1.2 The TruthfulQA Dataset
We chose to conduct our experiments using the TruthfulQA dataset [48] which meets these requirements. The dataset contains 817 distinct examples, which yields a standard error in the range of 1% to 2% for accuracy measures in the range of 90% to 50%. Each example contains a single query, and a variable number of correct and incorrect answers. And by considering each distinct $(q,a)$ pair as a demonstration for the purpose of building a context, the retriever is faced with similar demonstrations as multiple $(q,a)$ pairs share the same query (on average, the dataset contains 3.5 correct answers for each query).
4.1.3 Generating Demonstrations Pairs and Triplets
@@
  • $\mathcal{D}_{MC}$: this is the dataset as described in [48]. It contains 817 examples, each of which contains a variable number of correct and incorrect answers. The metrics MC1, MC2 and MC3, which can accept an arbitrary number of correct and incorrect answers as inputs, are calculated over this dataset.
@@
4.1.4 The Language Models
We conducted our experiments using four noteworthy LLMs: the smaller base text-completion model Mistral-7B-v0.1 (7B parameters), and the larger instruct-fine-tuned mixture of models Mixtral-8x22B-Instruct-v0.1 (141B parameters) from Mistral [50]; as well as a smaller chat-tuned model Llama-3-8B-chat (8B parameters) and a larger chat-tuned model Llama-3-70B-chat (70B parameters) from Llama  [^1] https://llama.meta.com/llama3/.
All four models are open-weights LLMs, meaning their internal parameters are publicly available for scrutiny and potential fine-tuning. These modern models stand out for achieving impressive performance on various tasks despite their relatively compact size. This efficiency makes it an attractive option for resource-constrained environments where deploying colossal models might not be feasible.
4.2 Implementation Details and Ablation Study
We implemented the Algorithm 1 and metrics from Section 3.2.1 in python. We computed BERT embeddings using the package sentence_transformers  [^2] https://sbert.net/, and implemented the retrieval algorithms in numpy. We queried all LLM models using the Together API  [^3] https://docs.together.ai/docs/inference-python.
We did not perform hyper-parameter tuning, and fixed the two parameters to $\lambda_{d}=0.75$ and $\lambda_{b}=0.95$ in all experiments. We fixed the amount of retrieved demonstrations to $k=6$, matching the number of few-shot examples from the fixed primer example from the TruthfulQA paper [48].
@@ 4.2 Implementation Details and Ablation Study
  • Fix: fixed primer examples [48]
  • Bias: Pure quality bias [12]
  • Rel: Pure semantic similarity [20] (KATE)
@@
  • Rel+Div: Semantic similarity plus diversity [26]
@@
4.3 Main Results
We present the experimental metrics for the 6 retrievers for the 4 different LLMs: in Table 2 and 2 for the Mistral models, and Table 4 and 4 for the Llama-3 models.
Table 1: Evaluation Metrics with Mistral-7B-v0.1
@@ 4.3 Main Results
Table 2: Evaluation Metrics with Mixtral-8x22B-Instruct-v0.1
Method DPO MC1 MC2 MC3
Fix -19.06 0.5202 0.3896 0.6799
Bias -27.30 0.4382 0.3096 0.5948
Rel -15.17 0.6193 0.5004 0.7616
Rel+Bias -14.77 0.6389 0.5080 0.7657
Rel+Div -12.67 0.6879 0.5181 0.8092
Rel+Div+Bias -13.29 0.6573 0.5071 0.7924
Table 3: Evaluation Metrics with Llama-3-8B-chat
@@
Table 4: Evaluation Metrics with Llama-3-70B-chat
Method DPO MC1 MC2 MC3
Fix -23.05 0.4382 0.3375 0.6184
Bias -20.41 0.4455 0.3303 0.6424
Rel -19.02 0.5483 0.4482 0.6958
Rel+Bias -19.09 0.5532 0.4495 0.7054
Rel+Div -13.93 0.6389 0.4834 0.7758
Rel+Div+Bias -13.72 0.6022 0.4621 0.7583
@@
The impact of few-shot learning is best seen by comparing the three MC metrics for Rel+Div for a smaller model against Fix for a larger model: the smaller models (7B and 8B parameters) enriched with ICL RAG are essentially matching or outperforming the bigger models (141B and 70B parameters) without ICL RAG.
The results consistently demonstrate that incorporating both relevance and diversity into the retrieval strategy leads to superior performance across all metrics and for both LLMs. For all models, and for all metrics, Rel+Div largely outperforms Rel. This finding reinforces the importance of not just retrieving relevant demonstrations but also ensuring a diverse set that maximizes the informative value for the LLM.
Interestingly, the impact of the low perplexity bias yields contrasting results. For both Mistral models, adding this bias results in a decline in performance on almost all metrics. Conversely, both Llama-3 models exhibit overall improvement with the low perplexity bias, in particular with the DPO metric. This intriguing observation suggests that LLM-dependent hyper-parameter tuning of $\lambda_{b}$ might be necessary to optimize retrieval strategies for specific models. Alternatively, the low perplexity bias itself may benefit from further refinement. Using an opposite intuition, we may argue that instead of prioritizing demonstrations the LLM already finds likely, introducing demonstrations that surprise the model the most could be beneficial for certain LLMs, potentially maximizing the learning impact of each demonstration. These findings open exciting new avenues for future research in ICL retrieval strategies, creating a parallel with novelty and serendipity concepts in recommender systems.
4.4 Calibrating Diversity using DPO
@@
Figure 1: Diversity Metric and DPO. Non-monotonous relationship between a diversity metric, the average cosine similarity between embedding pairs, and the quality metric DPO. Obtained by varying $\lambda_{d}$ with Llama-3-8B-chat and $k=6$.
5 Discussion: Real-World RAG Systems
@@
5.1 Balancing Performance and Efficiency
Retrieval latency is crucial at scale. Exhaustive, brute-force nearest neighbor search is computationally expensive and impractical. Instead, real-world systems leverage efficient indexing techniques and approximated kNN algorithms, as described in [23], to ensure fast retrieval times. This approach is essential for handling large datasets while maintaining responsiveness. To seamlessly integrate with existing retrieval engines and leverage their optimized search capabilities, a retrieval algorithm for RAG must ensure its data is stored in a format compatible with these engines. Commonly indexed data structures include text itself or low-dimensional dense vector embeddings. By adhering to these indexing practices, RAG systems can effectively leverage the power of existing retrieval engines and achieve fast, scalable retrieval of informative examples.
5.2 ICL RAG versus Fine-Tuning
@@
An interesting contrast between ICL and fine-tuning is highlighted in the paper [51]. The paper argues that fine-tuning can be more efficient than few-shot ICL in terms of cost and latency due to the super-linear increase in LLM latency with growing prompt sizes. However, this latency concern is less relevant with inference-throughput optimized LLM systems built with large GPU clusters, such as commonly used APIs. In these systems, the observed latency remains independent of the prompt size. From a latency perspective, adding ICL demonstrations can be considered free. Additionally, the paper suggests that ICL requires scanning through 100% of the demonstrations at query time. However this does not hold when employing real retrieval engines with indexing and approximate kNN, which significantly reduce the number of examples scanned during retrieval.
@@
5.3 Achieving State-of-the-Art Retrieval with Available Tools
@@
Due to the complexities of incorporating such rules directly within retrieval indexing algorithm [52], none of the solutions known to the authors from any of the above category implements diversity or quality biasing of the result. A common heuristic to mitigate this problem is to retrieve a larger set of candidate examples (e.g., double the desired number) and then apply diversity techniques like MMR on the retrieved candidates as a post-processing step. Quality biasing can be indirectly achieved by modifying the indexed embeddings themselves. For instance, reducing the norm of embeddings associated with low-quality content can nudge the retrieval algorithm towards higher-quality examples. An exact implementation in the context of cosine-similarity or dot-product relevance is to add an additional column storing the quality bias, and set the corresponding value to $1$ in the embedding of the query.
While vector search offers a powerful foundation for practical ICL retrieval, it often lacks native support for essential considerations like diversity or quality bias. These aspects are crucial for ensuring informative and effective retrieval of few-shot learning examples. Existing tools for recommendation engines, on the other hand, often excel in these areas. Recommendation engines natively incorporate rules that promote diversity by recommending a variety of items, or quality bias by prioritizing most popular products. Future research directions as well as practical systems for ICL retrieval could explore adapting or integrating these well-established techniques from recommender systems to further enhance the effectiveness and sophistication of few-shot learning through information retrieval. State-of-the-art ICL for real-world applications can be achieved by combining the strengths of vector search with the established "diversity-aware" retrieval approaches from recommender systems.
6 Conclusion